OpenStudy (anonymous):

Shaniqua starts a bank account with a $120 deposit. By the end of 12 months, she hopes to have saved $900. She knows that she can save $30 per month from her paycheck, but can earn some extra money by tutoring on the weekend. How much extra per month does Shaniqua need to save in order to meet her goal? Select the equation that best represents this situation and solution. A. 120 + 12(30 + x) = 900; $35 B. 120 + 12(30 + x) = 900; $420 C. 120 + 12(30) + x = 900; $35 D. 120 + 12(30) + x = 900; $420

OpenStudy (anonymous):

120 + 12(30 + x) = 900 120 + 360 + 12x = 900 480 + 12x = 900 12x = 900 - 480 12x = 420 x = 35 She would make $ 480 with just her paycheck savings plus the $ 120 starter money. Leaving her with $420 more to make to equal $ 900. If she makes $35 per month for her tutoring, she will have $ 900 in 12 months. Your answer is A :)